# Towards Effectively Detecting and Explaining Vulnerabilities Using Large Language Models

### v1

@@ 4.2. RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
  • Constructiveness: The explanation should provide actionable suggestions for code modification and remediation, offering guidance on how to fix the identified issue.
4.2.2. Experimental Results.
Table 6. Distribution (%) of Vulnerability Explanation Review Results (RQ2).
@@
We present the results of our dual evaluation process in Table 6, the suffix Gen. refers to the explanation generated by LLMVulExp and Ann. refers to the explanation annotated by GPT. All-Pos. refers to the results that meet requirements of all of the three metrics. Combining the results from manual and LLM evaluations, we have made the following observations:
(3) Constructiveness of Explanations: The explanatory generated by LLMVulExp generally includes actionable modification suggestions (i.e., 93.4% for SeVC and 80.5% for DiverseVul), indicating that training for vulnerability explanation can effectively help models provide practical remediation advice for developers.
(5) Potential of LLM in Automated Assessment: The automated assessment by the LLM is close to the capabilities of manual review. For SeVC and DiverseVul, the results of All-Pos. comparing manual review and automated review are 76.0% vs. 74.0% and 80.1% and 71.9%, respectively. This indicates that LLM has considerable potential in the evaluation of tasks related to the generation of vulnerability-related text. In the future, we can improve the accuracy of evaluations by using a dual verification process of manual and LLM automated reviews, while also reducing manual labor costs.
@@
Summary of RQ2: We find that LLMVulExp can generate explanations of the vulnerability with high Accuracy, Clarity, and Constructiveness. We also explore the potential of automated data annotation using LLMs to mitigate the manual effort.
4.3. RQ3: How do explanations affect the results of vulnerability detection?
@@
By considering the detection task as a step in the explanation task, we have conducted a comprehensive evaluation of the detection capabilities of the vulnerability explanation model in RQ1. This leads to an important research question: What is the impact of fine-tuning for vulnerability explanation on the detection capabilities of LLMs, compared to fine-tuning specialized detection models? In this RQ, we aim to experimentally explore the influence of explanatory information on the performance of vulnerability detection (i.e., comparing the performance of vulnerability detection between fine-tuned LLMs with and without the explanatory information).
@@
Experimental Setup. We conduct our research in the form of ablation studies under the settings of multi-type vulnerability detection scenario in SeVC (RQ1.3) and multi-label CWE detection scenario in DiverseVul (RQ1.4). We transform the vulnerability explanation fine-tuning task into a single vulnerability detection task by removing the explanatory information from the training data annotations and modifying the task description and instructions in the prompt. By doing so, we only require the LLMs to output the corresponding vulnerability types and then fine-tune the LLMs to obtain a vulnerability detection model. We compare the detection accuracy of the vulnerability detection model (referred to as Detector) with that of the vulnerability explanation model (referred to as Explainer).
Figure 5. Performance Comparison between Explainer and Detector.
@@
Experimental Results. We present the experimental results for comparison between the Explainer and the Detector on both datasets in Figure 5, the prefix of “Mi” and “Ma” refers to Micro and Macro. Based on the experimental results, incorporating the vulnerability explanation information into the fine-tuning process did not lead to a significant decline in detection performance. For example, the weightd precision on SeVC for Explainer and Detector are 80.1% and 81.1%, respectively, while 74.8% and 74.8% on DiverseVul. In fact, it even resulted in performance improvements under certain metrics on DiverseVul (e.g., the Explainer’s Macro-F1 score is 73.2%, which is nearly 2% higher than that of the Detector.).
@@
These observations indicate that the vulnerability explanation task can coexist with the vulnerability detection task without compromising the model’s accuracy in detecting vulnerabilities. The model’s detection capability still remains with an additional focus on explanation capabilities. Moreover, the automatically annotated vulnerability explanation data contains sufficient domain-specific knowledge, aiding the model in better understanding and identifying various vulnerability patterns. It suggests that enhancing a model’s explanation capabilities can be achieved without sacrificing detection performance. Furthermore, well-annotated explanatory data can improve the model’s overall understanding and performance in both vulnerability detection and explanation tasks.
@@
4.4. RQ4: How does the key code extraction impact detection performance?
In this RQ, we aim to investigate if the LLMs can identify the key code that might be prone to the vulnerability and further examine the code with a specific focus. To achieve this, we have conducted annotations of key code extraction on both the SeVC and DiverseVul datasets and fine-tuned the vulnerability explanation LLMs using the key code information.
Experimental Setup. Similar to the process of annotation discussed in Section 3, we use GPT-3.5 to extract the key code. The prompt templates are illustrated in Figure 4. To prevent label leakage, the code vulnerability type tags were not visible during the annotation process. For training data that does not output the original code statements, we will nullify its key code information. If it is test data, it will not be included in the evaluation. During fine-tuning and inference, we modify the prompts used, focusing on task description and instruction sections, to guide the model in focusing on the extracted key code for detecting the vulnerability.
Table 7. F1 Scores of Key Code Extraction on SeVC (RQ4)
@@ 4.4. RQ4: How does the key code extraction impact detection performance?
Experimental Results. We present the experimental results of utilizing Key Code Extraction on SeVC in Table 7 and DiverseVul in Figure 6. We find that there is a noticeable improvement for SeVC (i.e., 98.5% vs 92.4% for Single-Type and 83.8% vs 64.4% for Multi-Type). For DiverseVul, there is a back-and-forth trend comparing the results with key code and without key code. The potential reason might be that the longer and more complex code length makes it difficult to effectively extract the key code on DiverseVul. The results indicate that this enhancement scheme can considerably improve the detection accuracy of the model in different vulnerability explanation tasks. This demonstrates the importance of supplementing code semantic information for the fine-tuning of large vulnerability models and the feasibility of semantic information extraction based on large models. On the one hand, by extracting key portions of the code that are more likely to contain vulnerabilities, the model can concentrate on the most pertinent information, leading to more accurate detection results. On the other hand, key code extraction helps eliminate irrelevant parts of the code, reduce noise, and improve the signal-to-noise ratio. This makes it easier for the model to learn and detect patterns associated with vulnerabilities.
Figure 6. Performances of Utilizing Key Code Extraction on DiverseVul Dataset(RQ4), the numbers in x-axis represent for the CWE IDs.
Summary of RQ4: By guiding the LLMs to focus on key code, the performance of vulnerability detection can be improved by a large margin (e.g., 64.4% -¿ 83.8% for Multi-Type detection on SeVC).
5. Discussion
@@
Implication 1: Implications for Broader Vulnerability-Related Tasks. Our study reveals that integrating the task of vulnerability explanation with detection does not compromise the model’s performance. This finding suggests we should consider a wider range of vulnerability-related tasks beyond mere detection when fine-tuning LLMs. By doing so, we can potentially enhance the model’s overall understanding and its ability to provide more contextually rich and actionable insights. For instance, tasks such as vulnerability impact assessment, prioritization based on severity, or even automated patch generation could be integrated into the training regime. This holistic approach could lead to the development of more sophisticated tools that not only identify vulnerabilities but also assist in managing and mitigating associated risks.
Implication 2: Effectiveness of LLMVulExp Using Different LLMs. In this paper, we use CodeLlama-13B-Instruct as the primary model to conduct the experiments. To investigate the effectiveness of our framework using different LLMs, we utilize two additional LLMs (i.e., CodeLlama-7B-Instruct and Llama3-8B-Instruct) to conduct the experiments following the setting of RQ1.3. Figure 7 shows the results of Multi-Type vulnerability detection using LLMs with different sizes. We find that LLMVulExp can also achieve effective performance using the additional LLMs. Particularly, the performance between CodeLlama-13B-Instruct and CodeLlama-7B-Instruct is similar and outperforms Llama3-8B-Instruct. Overall, LLMVulExp presents its effectiveness using different LLMs with various sizes.
Figure 7. Performance of Multi-Type vulnerability detection using LLMs with different sizes.
@@ 6. Threats to Validity
Construct Validity. We evaluated the explanatory capability using three proposed metrics: Accuracy, Clarity, and Constructiveness. These metrics characterize three aspects of effective vulnerability explanations but may not comprehensively cover all features of vulnerability explanations. More comprehensive and refined evaluation metrics for vulnerability explanations require further research. In this paper, we employed both manual review and automated review using LLMs. To mitigate the bias, two authors independently annotated the explanations and resolved any discrepancies until a consensus. The value of Cohen’s Kappa (McHugh, 2012) in this process is 0.76, which indicates a substantial agreement.
@@
Internal Validity. Due to the high computational cost of fine-tuning and inference evaluation of LLMs, we were unable to conduct experiments under different data splits and randomization states. The randomness in our experiments (e.g., data splitting, explanation annotation, and fine-tuning process) may affect the results. However, we conducted at least two or more trials to validate the stability of our conclusions for each experimental setup. The number of hyperparameter trials for LoRa configuration and the LLM generation method was relatively limited. Additionally, due to GPU constraints, we did not attempt to fine-tune larger models, such as the 34B version. Although these limitations may have resulted in alternative findings, they do not fundamentally undermine our primary objective: to explore the potential of conducting vulnerability detection and explanation tasks using LLMs.
@@
External Validity. We chose the advanced open-source code LLM CodeLlama, based on the Llama2 architecture, as our primary model for this research. We tested different versions, including 7B and 13B, and evaluated the performance of the recently released general large language model Llama3 on vulnerability detection and explanation tasks. However, there are many different architectures of general-purpose open-source LLMs and code models, and our experiments were limited to a few models. We conducted experiments using two differently configured open-source project-based vulnerability datasets. While these datasets contain a lot of vulnerable code, they differ from real-world development environments. Therefore, our findings may not apply to other datasets or industrial settings.
7. Conclusion
@@
In this paper, we propose LLMVulExp, a framework that aims that detecting and explaining vulnerabilities using LLMs. The results underscore the potential of LLMs in advancing vulnerability detection and explanation in software security. Our research provides valuable insights into the fine-tuning of LLMs for vulnerability detection and explanation. It suggests that a multi-task learning approach could be beneficial, addressing the data volume bottleneck is crucial for practical applications, and the quality of annotations is paramount for the success of fine-tuning frameworks. By considering these insights, future work can aim to develop more capable and efficient models that contribute to the field of software security.

### v2

@@ IV-B1 Experimental Setup
  • Actionability: The explanation should provide actionable suggestions for code modification and remediation, offering guidance on how to fix the identified issue.
IV-B2 Experimental Results
We present the results of our dual evaluation process in Table VIII, the suffix Gen. refers to the explanation generated by LLMVulExp and Ann. refers to the explanation annotated by GPT-3.5. All-Pos. refers to the results that meet the requirements of all of the three metrics. Combining the results from manual and LLM evaluations, we have made the following observations:
(3) Actionability of Explanations: The explanatory generated by LLMVulExp generally includes actionable modification suggestions (i.e., 93.4% for SeVC and 80.5% for DiverseVul), indicating that training for vulnerability explanation can effectively help models provide practical remediation advice for developers.
(5) Potential of LLM in Automated Assessment: The automated assessment by the LLM is close to the capabilities of manual review. For SeVC and DiverseVul, the results of All-Pos. comparing manual review and automated review are 76.0% vs. 74.0% and 80.1% vs. 71.9%, respectively. This indicates that LLM has considerable potential in the evaluation of tasks related to the generation of vulnerability-related text. In the future, we can improve the accuracy of evaluations by using a dual verification process of manual and LLM automated reviews, while also reducing manual labor costs.
TABLE VIII: Distribution (%) of Vulnerability Explanation Review Results (RQ2).
@@ IV-B2 Experimental Results
1int i; 2for(i = 0; i < 1; i++) 3twoIntsStruct * dataBuffer = (twoIntsStruct *) ALLOCA(100 * sizeof(twoIntsStruct)); 4size_t i; 5for (i = 0; i < 100; i++) 6 dataBuffer[i].intOne = 1; 7 dataBuffer[i].intTwo = 1; 8data = dataBuffer; 9printStructLine(&data[0]); 10 11void printStructLine(const twoIntsStruct * structTwoIntsStruct) { 12 printf($d\n$", structTwoIntsStruct->intOne, structTwoIntsStruct->intTwo); 13} 14delete [] data;
{mdframed}[style=mystyle] [type] pointer
[location] The line ”delete [] data” has a pointer-related issue.
[explanation]
(Analysis:) The issue in this code is related to the ”data” pointer, which is being improperly freed using the ”delete” operator. The ”data” pointer is assigned the address of a memory block allocated on the stack using the ”ALLOCA” function, which allocates memory on the stack instead of the heap. The ”delete” operator is used to free memory allocated on the heap using the ”new” operator, and using it to free memory allocated on the stack may cause undefined behavior or crashes.
(Suggestion:) To fix this issue, the program should use the ”free” function to free memory allocated on the stack, or use the ”new” operator to allocate memory on the heap instead of the ”ALLOCA” function. Additionally, the program should ensure that the ”data” pointer points to a valid memory location before it is used or accessed, and should add proper error handling and validation to ensure that the ”data” pointer behaves correctly and safely in all cases.
TABLE IX: Example of Generated Explanation on SeVC.
IV-B3 Case Study
To more intuitively illustrate the importance of evaluating the quality of vulnerability explanations, we analyze an example generated by our model. As shown in Figure IV-B2, this vulnerability code is sourced from SeVC and pertains to pointer-type vulnerabilities, demonstrating incorrect memory deallocation in C language using the delete operator. The explanation adheres to our output paradigm, generating vulnerability detection, location, and detailed information.
The provided explanation correctly identifies the core issue: the data pointer is allocated on the stack using the ALLOCA function but is improperly freed using the delete operator. The explanation points out that the delete operator should only be used to free memory allocated on the heap via the new operator. Using it to free stack-allocated memory can lead to undefined behavior and potential crashes, meeting the accuracy requirement. Additionally, the explanation offers recommendations for addressing this vulnerability, such as using the free function to release stack-allocated memory or using the new operator for heap allocation, and ensuring pointer safety. This satisfies the actionability requirement. The explanation is clear, comprehensive, logically coherent, and easy to read, meeting the clarity requirement.
During the review, we also observe some instances with insufficient quality of explanations. Issues include failure to identify the main vulnerability, vague location, verbose expression, and lack of remediation suggestions. Hallucinations in explanation generation are relatively rare but primarily involved memorized descriptions of similar code samples from the training set and location confusion. Overall, the fine-tuned model demonstrates a certain capability for generating vulnerability explanations.
@@ IV-B3 Case Study
Summary of RQ2: We find that LLMVulExp can generate explanations of the vulnerability with high Accuracy, Clarity, and Actionability. We also explore the potential of automated data annotation using LLMs to mitigate the manual effort.
IV-C RQ3: How do explanations affect the results of vulnerability detection?
@@
Considering the detection task as a step in the explanation task, we have conducted a comprehensive evaluation of the detection capabilities of the vulnerability explanation model in RQ1. This leads to an important research question: What is the impact of fine-tuning for vulnerability explanation on the detection capabilities of LLMs compared to fine-tuning specialized detection models? In this RQ, we aim to experimentally explore the influence of explanatory information on the performance of vulnerability detection by comparing the performance of fine-tuned LLMs with and without explanatory information.
@@
Experimental Setup. We conduct our research as ablation studies under the settings of multi-type vulnerability detection scenarios in SeVC (RQ1.3) and multi-label CWE detection scenarios in DiverseVul (RQ1.4). We transform the vulnerability explanation fine-tuning task into a single vulnerability detection task by removing the explanatory information from the training data annotations and modifying the task description and instructions in the prompt. By doing so, we only require the LLMs to output the corresponding vulnerability types and then fine-tune the LLMs to obtain a vulnerability detection model. We compare the detection accuracy of the vulnerability detection model (referred to as Detector) with that of the vulnerability explanation model (referred to as Explainer).
Figure 4: Performance Comparison between Explainer and Detector.
@@
Experimental Results. We present the experimental results for comparison between the Explainer and the Detector on both datasets in Figure 4. The prefix of “Mi” and “Ma” refers to Micro and Macro, respectively. Based on the experimental results, incorporating vulnerability explanation information into the fine-tuning process did not lead to a significant decline in detection performance. For example, the weighted precision on SeVC for the Explainer and Detector are 80.1% and 81.1%, respectively, while 74.8% for both on DiverseVul. In fact, it even resulted in performance improvements under certain metrics on DiverseVul (e.g., the Explainer’s Macro-F1 score is 73.2%, which is nearly 2% higher than that of the Detector.).
@@
These observations indicate that the vulnerability explanation task can coexist with the vulnerability detection task without compromising the model’s accuracy in detecting vulnerabilities. The model’s detection capability remains intact with an additional focus on explanation capabilities. Moreover, the automatically annotated vulnerability explanation data contains sufficient domain-specific knowledge, aiding the model in better understanding and identifying various vulnerability patterns. This suggests that enhancing a model’s explanation capabilities can be achieved without sacrificing detection performance. Furthermore, well-annotated explanatory data can improve the model’s overall understanding and performance in both vulnerability detection and explanation tasks.
@@
IV-D RQ4: How does the key code extraction impact detection performance?
In this RQ, we aim to investigate if the LLMs can identify the key code that might be prone to vulnerability and further examine the code with a specific focus. To achieve this, we have conducted annotations of key code extraction on both the SeVC and DiverseVul datasets and fine-tuned the vulnerability explanation LLMs using the key code information.
Experimental Setup. Similar to the process of annotation discussed in Section III, we use GPT-3.5 to extract the key code. To prevent label leakage, the code vulnerability type tags were not visible during the annotation process. For training data that does not output the original code statements, we will nullify its key code information. If it is test data, it will not be included in the evaluation. During fine-tuning and inference, we modify the prompts used, focusing on task description and instruction sections, to guide the model in focusing on the extracted key code for detecting the vulnerability.
TABLE X: F1 Scores of Key Code Extraction on SeVC (RQ4).
@@ IV-D RQ4: How does the key code extraction impact detection performance?
Experimental Results. We present the experimental results of utilizing Key Code Extraction on SeVC in Table X and DiverseVul in Figure 5. We find that there is a noticeable improvement for SeVC (i.e., 98.5% vs 92.4% for Single-Type and 83.8% vs 64.4% for Multi-Type). For DiverseVul, there is a back-and-forth trend comparing the results with key code and without key code. The potential reason might be that the longer and more complex code length makes it difficult to effectively extract the key code on DiverseVul. The results indicate that this enhancement scheme can considerably improve the detection accuracy of the model in different vulnerability explanation tasks. This demonstrates the importance of supplementing code semantic information for the fine-tuning of large vulnerability models and the feasibility of semantic information extraction based on large models. On the one hand, by extracting key portions of the code that are more likely to contain vulnerabilities, the model can concentrate on the most pertinent information, leading to more accurate detection results. On the other hand, key code extraction helps eliminate irrelevant parts of the code, reduce noise, and improve the signal-to-noise ratio. This makes it easier for the model to learn and detect patterns associated with vulnerabilities. This method does not rely on manual feature extraction, making it broadly applicable. Our experimental results preliminarily demonstrate the feasibility of key code extraction through LLMs.
Figure 5: Performances of Utilizing Key Code Extraction on DiverseVul Dataset(RQ4), the numbers in x-axis represent the CWE IDs.
Summary of RQ4: By guiding the LLMs to focus on key code, the performance of vulnerability detection can be improved by a large margin (e.g., from 64.4% to 83.8% for Multi-Type detection on SeVC).
V Discussion
@@
Implication 1: Broader Applications for Vulnerability-Related Tasks. Our study reveals that integrating vulnerability explanation with detection does not compromise the model’s performance. This finding suggests that we should consider a wider range of vulnerability-related tasks beyond mere detection when fine-tuning LLMs. By doing so, we can potentially enhance the model’s overall understanding and its ability to provide more contextually rich and actionable insights. For instance, tasks such as vulnerability impact assessment, prioritization based on severity, or even automated patch generation could be integrated into the training regime. This holistic approach could lead to the development of more sophisticated tools that not only identify vulnerabilities but also assist in managing and mitigating associated risks.
Implication 2: Effectiveness of LLMVulExp Using Different LLMs. In this paper, we use CodeLlama-13B-Instruct as the primary model to conduct the experiments. To investigate the effectiveness of our framework using different LLMs, we utilize two additional LLMs, CodeLlama-7B-Instruct and Llama3-8B-Instruct, to conduct the experiments following the setting of RQ1.3. Figure 6 shows the results of multi-type vulnerability detection using LLMs of different sizes. We find that LLMVulExp can achieve effective performance with these additional LLMs. Notably, the performance of CodeLlama-13B-Instruct and CodeLlama-7B-Instruct is similar and outperforms Llama3-8B-Instruct. Overall, LLMVulExp demonstrates its effectiveness using different LLMs of various sizes.
Figure 6: Performance of Multi-Type vulnerability detection using LLMs with different sizes.
@@ VI Threats to Validity
Construct Validity. We evaluated the explanatory capability using three proposed metrics: Accuracy, Clarity, and Actionability. While these metrics effectively characterize important aspects of vulnerability explanations, they may not comprehensively cover all features. More comprehensive and refined evaluation metrics for vulnerability explanations require further research. To mitigate potential bias, we employed both manual and automated reviews using LLMs. Two authors independently annotated the explanations and resolved any discrepancies until a consensus was reached. The value of Cohen’s Kappa [54] in this process was 0.76, indicating substantial agreement. This rigorous approach helps ensure the reliability and validity of our evaluation.
@@
Internal Validity. Due to the high computational cost of fine-tuning and inference evaluation of LLMs, we were unable to conduct experiments under different data splits and randomization states. The randomness in our experiments (e.g., data splitting, explanation annotation, and fine-tuning process) may affect the results. However, we conducted multiple trials to validate the stability of our conclusions for each experimental setup. Although the number of hyperparameter trials for LoRa configuration and the LLM generation method was relatively limited, and due to GPU constraints, we did not attempt to fine-tune larger models, such as the 34B version, we ensured that these limitations did not fundamentally undermine our primary objective: to explore the potential of conducting vulnerability detection and explanation tasks using LLMs.
@@
External Validity. We chose the advanced open-source code LLM CodeLlama, based on the Llama2 architecture, as our primary model for this research. We tested different versions, including 7B and 13B, and evaluated the performance of the recently released general large language model Llama3 on vulnerability detection and explanation tasks. However, there are many different architectures of general-purpose open-source LLMs and code models, and our experiments were limited to a few models. Additionally, we conducted experiments using two differently configured open-source project-based vulnerability datasets. Although these datasets include a relatively large amount of vulnerable code, there are still differences compared to code in real development environments. Consequently, our experimental conclusions have an inherent risk of not generalizing to other vulnerability datasets or industrial environments. Nonetheless, our careful selection and thorough evaluation process aimed to ensure that our findings remain relevant and insightful within the scope of our study.
VII Conclusion
@@
In this paper, we propose LLMVulExp, a framework designed to detect and explain vulnerabilities using LLMs. The results underscore the potential of LLMs in advancing vulnerability detection and explanation in software security. Our research provides valuable insights into the fine-tuning of LLMs for vulnerability detection and explanation. It highlights the importance of addressing the data volume bottleneck for training vulnerability LLMs in software development. Furthermore, the quality of annotations is paramount for the success of fine-tuning frameworks. Considering these insights, future work can aim to develop more capable and efficient models that significantly contribute to the field of software security.